# Hyperparameter Optimization in Machine Learning

### v1

@@ 2.5 Tackling the search manually
Yet manual search can quickly become exhausting. Experienced practitioners may intuitively narrow the search, but novices face a dizzying array of choices. While flexible, tweaking hyperparameters without a systematic approach risks getting lost in an endless maze of configurations. Furthermore, with each trial, insights and analysis of results can guide exploration. While potentially beneficial to speed up the process, this risks overfitting by inadvertently exploiting patterns in the validation data. Manual search also suffers from poor reproducibility. The winding path through the hyperparameter space exists only in the searcher’s mind, impossible to retrace.
@@
We start with grid search, an approach that consists in dividing the search space into a grid of values and querying the response function on each of these values. Next, random search swaps the grid in favour of a fixed, simple distribution over the search space, drawing configurations from it until a budget is exhausted. Finally, we move to quasi-random search strategies, approaches rooted in low-discrepancy sequence generation, in which the simple distributions of random search are replaced by more complex schemes that seek to explore the space more evenly. While grid search is a traditional and classic technique in HPO, consensus has grown in the field that random search should be considered as the baseline method for HPO [bergstra_random_2012]. Quasi-random search may offers benefits over random search, although to date we are not aware of any theoretical result supporting this.
@@ 3.1 Grid Search
The search grid is then given by all the combinations on the Cartesian product generated by these $m$ sets, namely $|\upsilon|=\prod_{i=1}^{m}n_{i}$. The learning algorithm can be executed sequentially or in parallel on the resulting set of points. The hyperparameter configuration that minimizes the objective function $f$ over $\upsilon$ is returned as $\widehat{\lambda}$; that is
$\widehat{\lambda}\in\operatorname*{arg\,min}_{\lambda\in\upsilon}f(\lambda).$ (3.1)
@@ 3.2 Random Search
Perhaps conceptually the simplest procedure for HPO is random search (RS), which was already considered in the 50’ as a method for global optimization [brooks1958discussion]. RS consists in repeatedly drawing hyperparameter configurations from a predefined distribution $\upsilon$ over $\Lambda$ and it is readily parallelized. The incumbent solution after $K$ trials is given by
$\widehat{\lambda}_{K}=\operatorname*{arg\,min}_{(\lambda_{1},\dots\lambda_{k})\sim\upsilon^{K}}f(\lambda_{i}),$ (3.2)
@@
where the trial points $\lambda_{i}$ are drawn i.i.d. from $\upsilon.$ While in GS the total number of trial points is predetermined, one may keep on executing RS indefinitely. Termination conditions might involve approximate convergence checks, like lack of improvements after a certain amount of trials, or, much more frequently in HPO, after a budget is exhausted (see Section 2.4), e.g., one may run RS for 12 hours.
@@
Even if in principle $\upsilon$ could be any distribution, when referring to random search for hyperparameter optimization, one typically implies the use of simple distributions, such as uniform or log-uniform, that factorize independently along the configuration components, i.e. $\upsilon=\prod_{j=1}^{m}\upsilon_{i}$. For instance, if $\lambda$ contains, in order, a categorical choice between $K$ different neural architectures, the learning rate for stochastic gradient descent and a global dropout factor for regularization then $\upsilon$ could be set as
$\upsilon=\mathcal{U}\{1,K\}\times\mathrm{Log}\,\mathcal{U}(10^{-5},10^{-1})\times\mathcal{U}(0.5,1).$
where $\mathcal{U}\{\cdot\}$ and $\mathcal{U}(\cdot)$ are the discrete and continuous uniform distributions and $\mathrm{Log}\,\mathcal{U}(\cdot)$ is the log-uniform (or reciprocal) distribution on the indicated intervals.
On the efficiency of random search (compared to grid search).
@@
In an influential article, bergstra_random_2012 advocate for using RS (instead of manual search or grid search) as the “default” baseline method for HPO. They argue that while RS retains the advantages of GS, such as reproducibility, conceptual simplicity, very low implementation overhead and ease of parallelization, it is also more efficient in higher dimensional spaces, especially when the response function has a low effective dimensionality. The authors informally define this condition as when a subset of the hyperparameters accounts for most of the variation of the response function. For instance, for a bi-dimensional hyperparameter it could be that $f(\lambda_{1},\lambda_{2})\approx g(\lambda_{1})$.
@@
One way to reason about the success of RS is to assume that there is a sub-region $\Lambda^{*}\subset\Lambda$ that leads to a positive outcome of the experiment, i.e., where $f$ is sufficiently small, and compare the volume (under random sampling from $\upsilon$) of such region with that of the full search space [brooks1958discussion]. We can model the probability of success of random search after $K$ draws as follows:
$\mathbb{P}_{\upsilon}\left(\widehat{\lambda}_{K}\in\Lambda^{*}\right)=1-\left(1-\mathbb{P}_{\upsilon}(\lambda\in\Lambda_{*})\right)^{K},$ (3.3)
@@
where $\mathbb{P}_{\upsilon}$ denotes the probability of sampling from $\upsilon$, typically uniform or log-uniform. Eq. (3.3) says that the minimum number of trials needed to assure a positive outcome with a given probability, say $95\%$, depends only on the probability of hitting $\Lambda_{*}$ under $\upsilon$. This can be indeed independent of the search space dimension if the response function has a low effective dimensionality, argument that may explain the success of RS as a baseline in seemingly complex settings such as neural architecture search [li2019random], where repeated experimentation on a few well-known benchmark datasets has led to the development of well-tuned search spaces for neural architectures, where the probability of hitting a favorable configuration by random sampling is relatively high.
@@
However, for a general search space, we should expect $\mathbb{P}_{\upsilon}(\lambda\in\Lambda_{*})$ to decay exponentially in some $m^{\prime}\leq m$. For concreteness, consider, for instance, the case where $\Lambda=[0,1]^{m}$ is the $m$-dimensional unit hypercube and $\Lambda^{*}=[0,1/2]^{m}$ is the $m$-dimensional $1/2$ hypercube. If $\upsilon=\mathcal{U}(\Lambda)$ is the uniform distribution over $\Lambda$, then $\mathbb{P}_{\upsilon}(\lambda\in\Lambda_{*})=2^{-m}$. Consequently, the number of trials needed for having a positive outcome with at least $95\%$ probability is $K\geq\log(0.05)/\log(1-2^{m})$, which grows exponentially in $m$. For $m=3$ one already needs $K\geq 23$ draws to achieve a $95\%$ probability of success. On the other hand, in a low effective dimensionality setting we could have that $\Lambda^{*}=[0,1/2]^{m^{\prime}}\times[0,1]^{m^{\prime\prime}}$ with $m^{\prime}+m^{\prime\prime}=m$. It is immediate to see that $K$ grows with $m^{\prime}$ only, illustrating the fact that RS is insensitive to to the number of inactive dimensions, covering a larger number of distinct values compared to GS.
An additional advantage of random search over grid search is that RS allows for a more natural treatment of a larger class of hyperparameters, not requiring discretization of continuous search spaces and allowing for distributions with non-compact supports. However, the randomized nature of this HPO technique means that “unlucky” runs may end up with entire regions of $\Lambda$ being under-explored or completely neglected. As a consequence, a single execution of random search may not necessarily provide sufficient information to meaningfully modify $\upsilon$ for another iteration of random search, should it be needed. Continuing with our illustrative examples, say we divide the unit hypercube in $2^{m}$ 1/2 hypercubes and we wish to have one trial per hypercube (anywhere). While GS would need exactly $2^{m}$ points to achieve this, the probability under uniform sampling that RS hits all the hypercubes in $2^{m}$ draws is $\prod_{j=1}^{2^{m}-1}(1-j/2^{m})$. For example, for $m=3$ such probability is already $3/32$, meaning that, in expectation, RS with $8$ draws covers all the hypercubes only once every ten runs.
@@ 3.3 Quasi-random search
One basic requirement for a stochastic generator of hyperparameter configuration is consistency: we would like that $\lim_{k\to\infty}\min_{i\in[k]}f(\lambda_{i})=f(\lambda^{*})$. RS with a uniform distribution is consistent. However, unlike GS, as we have discussed above, for a fixed budget of draws RS may generate a set of configurations that misses large portions of the search space. Intuitively, good sequences should cover evenly the search space. Also their projections onto subspaces of $\Lambda$ should be well spread, in order to speed up convergence when optimizing functions with low effective dimensionality; a property that GS is missing. Formalizations of these desiderata have been developed in the context of numerical integration and global optimization via the concepts of low dispersion and low discrepancy sequences [see e.g. sobol1979systematic, niederreiter1988low, dick2010digital]. bousquet2017critical study the usage of such quasi-random sequences in HPO, finding that they consistently outperform random search (with a uniform distribution) on a series of deep learning tasks. The resulting method is conceptually very similar to both GS and RS, being highly parallelizable and easy-to-implement, except that the set of hyperparameters $\lambda_{1},\dots,\lambda_{k}$ are generated according to more complex procedures. We review a few possible choices below, assuming that the search space is an $m$-dimensional unit hypercube for simplicity. One can then derive samples for a variety of distributions with appropriate reparameterizations [devroye2006nonuniform].
Latin hypercube sampling (LHS) [mckay1979comparison].
Choose a sequence length $n>1$, which determines the number of hyperparameters generated by a run of the sampling process. Divide each $[0,1]$ interval of the unit hypercube into $n>1$ parts, denoting by $I_{s}=[(s-1)/n,s/n]$ the resulting intervals. This gives rise to $n^{m}$ hypercubes of volume $n^{-m}$ each. Then generate $m$ random permutations of $[n]$, denoted by $\{\sigma_{i}\}_{i=1}^{m}$, $\sigma_{i}:[n]\to[n]$. For $k\in[n]$, sample hyperparameter configurations as follows:
@@
and $\mathcal{U}(H_{k})$ is the uniform distribution on $H_{k}$. This generates $n$ trial points; one may then generate further sequences of $n$ configurations by repeating the process. Latin hypercube sampling ensures that the projection of the sequence on each axis will place exactly one sample per interval.
Sobol’ sequences [sobol1979systematic].
Sobol’ sequences are a collection of separate one-dimensional sequences, one for each axis: when constructing a $m+m^{\prime}$ dimensional sequence one can reuse the $m$-dimensional one for the first coordinates.
Let $j\in[m]$ be an axis, and let $\mathbb{F}_{2}=\{0,1\}$ with integer operations modulo $2$. For each integer $k$ (the iterate) let $k=\sum_{l}a_{l}(k)2^{j}$ be the binary expansion of $k$, where $a_{l}(k)\in\mathbb{F}_{2}$ are binary coefficients (e.g., for $k=3$, these would be $a_{0}(3)=1$, $a_{1}(3)=1$ and the rest 00’s). Denote by $a(k)$ the resulting vector. Let $C_{j}$ be a specific binary (infinite) “direction matrix” for axis $j$. Then the $j$-th coordinate of the $k$-th iterate of a Sobol’ sequence can be written as
@@
where the matrix-vector multiplication $C^{j}\cdot a(k)$ is in $\mathbb{F}_{2}$ (i.e., addition is modulo 2). The direction matrices $C^{j}$ are constructed starting from primitive polynomials of $\mathbb{F}_{2}$ and direction numbers, which must be carefully chosen to ensure good coverage of the sequences. We refer the reader to [Lemieux:1338331, Chapter 5] for details.
Note that, given the directions matrices, Sobol’ sequences are deterministic. It is common practice to obtain stochastic variants of such sequences by adding random shifts, i.e. $\lambda^{\prime}_{k}=\lambda_{k}+u$ (modulo the unit hypercube) with $u\sim\mathcal{U}([0,1]$ and by “scrambling” the binary representation vector $a(k)$, that is randomly permuting its entries [owen1995randomly].
Other schemes.
@@
Figure 3.1: Non-adaptive sampling methods on the unit square. The main plot shows a sequence of samples generated by different methods. For all methods except Latin Hypercube Sampling (LHS), the initial 4 samples are represented by darker and larger circles, followed by 5 additional samples depicted as lighter and smaller circles. For LHS, a sequence of 9 samples is directly shown, as this method requires resampling for each selected number of draws. The points on the top and right of the main plot are projections of the samples onto the respective axes. We can observe that the grid sampling method covers only two (with 4 draws) and three (with 9 draws) distinct points on the axes projections, indicating poor coverage. The first four draws of uniform random sampling illustrate the “unlucky run” phenomenon where the points concentrate in a small region of the square. Even after 9 draws, random sampling leaves the first quadrant unvisited. In contrast, Latin Hypercube Sampling and Sobol’ sequences (with random shifts and scrambling) provide more even coverage of the space.
@@
The techniques described in the previous chapter do not make any attempt to model the response function $f(\lambda)$. Except for manual search, the information retrieved on the previous runs of the learning algorithm does not influence the search. In contrast, model-based (MB) techniques construct an explicit surrogate model for $f(\lambda)$ – or for closely related quantities – based on the information collected from previous evaluations. We will denote such model by
@@
The landscape of model-based methods is strongly dominated by techniques falling under the umbrella of Bayesian optimization (BO), which we describe in more details in the next section. In BO the surrogate model is a probabilistic mapping that also captures the uncertainty of the estimation. Nonetheless, some recent works go in different directions. ilievski2017efficient use a deterministic radial basis function model for $\hat{f}$ which they interpolate on observations of $f$, while hazan2017hyperparameter focus on Boolean hyperparameters and use a sparse Fourier basis model for $\hat{f}$, in conjunction with heuristics to estimate the most sensitive hyperparameter entries. They then use a base global optimizer for performing HPO on a restriction of the original hyperparameter space. The works by lorraine2018stochastic and mackay2019self, linked to gradient-based HPO may also be interpreted as an attempt to learn a surrogate model of the entire learning algorithm (applied to a specific dataset), whereby the output space of $\hat{f}$ coincides with the hypothesis space of $A$ itself, that is $\hat{\mathcal{Y}}=\mathcal{H}$ in (4.1).
Regardless of the particular implementation, one critical feature of surrogate models is that they should be much cheaper to evaluate than $f$ itself (we recall that evaluating $f$ requires executing the learning algorithm $A$, which may be very expensive). This typically allows model-based techniques to compute $\hat{f}$ on several values at each step, e.g., to seek for extrema. Another critical feature of surrogate models is the ability to provide uncertainty estimations for evaluations of $f$. This allow the optimization process to balance exploration, i.e., collecting information in areas of high uncertainty to improve the model, and exploitation, i.e., leveraging the learned model to sample hyperparameters close to the current minimizer.

### v2

@@ 2.5 Tackling the search manually
Yet manual search can quickly become exhausting. Experienced practitioners may intuitively narrow the search, but novices face a dizzying array of choices. While flexible, tweaking hyperparameters without a systematic approach risks getting lost in an endless maze of configurations. Furthermore, with each trial, insights and analysis of results can guide exploration. While potentially beneficial to speed up the process, this risks overfitting by inadvertently exploiting patterns in the validation data. Manual search also suffers from poor reproducibility, as it is hard to track the steps an user followed to reach some final configuration.
@@
We start with grid search, an approach that consists in dividing the search space into a grid of values and querying the response function on each of these values. Next, random search swaps the grid in favour of a fixed, simple distribution over the search space, drawing configurations from such distribution until a budget is exhausted. Finally, we move to quasi-random search strategies, approaches rooted in low-discrepancy sequence generation, in which the simple distributions of random search are replaced by more complex schemes that seek to explore the space more evenly. While grid search is a traditional and classic technique in HPO, consensus has grown in the field that random search should be considered as the baseline method for HPO [bergstra_random_2012]. Quasi-random search may offers benefits over random search, although to date we are not aware of any theoretical result supporting this.
@@ 3.1 Grid Search
The search grid is then given by all the combinations on the Cartesian product generated by these $m$ sets, namely $|\upsilon|=\prod_{i=1}^{m}n_{i}$. The learning algorithm can be executed sequentially or in parallel on the resulting set of points. The hyperparameter configuration that minimizes the objective function $f$ over $\upsilon$ is returned as $\hat{\lambda}$; that is
$\hat{\lambda}\in\operatorname*{arg\,min}_{\lambda\in\upsilon}f(\lambda).$ (3.1)
@@
The top row of Figure 3.1 depicts equally distanced grid “samples” with 4, 9 and 16 points, on the unit square.
@@ 3.2 Random Search
Perhaps conceptually the simplest procedure for HPO, random search (RS) appears already considered in the 50’ as a method for global optimization [brooks1958discussion]. RS consists in repeatedly drawing hyperparameter configurations from a predefined distribution $\upsilon$ over $\Lambda$ and it is readily parallelized. The incumbent solution after $K$ trials is given by
$\hat{\lambda}_{K}=\operatorname*{arg\,min}_{(\lambda_{1},\dots\lambda_{k})\sim\upsilon^{K}}f(\lambda_{i}),$ (3.2)
@@
where the trial points $\lambda_{i}$ are drawn i.i.d. from $\upsilon.$ While in GS the total number of trial points is predetermined, one may keep on executing RS indefinitely. Termination conditions might involve approximate convergence checks, like lack of improvements after a certain amount of trials, or, much more frequently in HPO, after a budget is exhausted (see Section 2.4). For instance, one may run RS for 12 hours.
@@
Even if in principle $\upsilon$ could be any distribution, when referring to random search for hyperparameter optimization, one typically implies the use of simple distributions (uniform or log-uniform distributions) that factorize independently along the configuration components, i.e. $\upsilon=\prod_{j=1}^{m}\upsilon_{i}$. For instance, if $\lambda$ contains, in order, a categorical choice between $K$ different neural architectures, the learning rate for stochastic gradient descent and a global dropout factor for regularization then $\upsilon$ could be set as
$\upsilon=\mathcal{U}\{[k]\}\times\mathrm{Log}\,\mathcal{U}(10^{-5},10^{-1})\times\mathcal{U}(0,0.5).$
where $\mathcal{U}\{\cdot\}$ and $\mathcal{U}(\cdot)$ are the discrete and continuous uniform distributions and $\mathrm{Log}\,\mathcal{U}(\cdot)$ is the log-uniform (or reciprocal) distribution on the indicated intervals. Using a distribution rather than a grid allows for a more natural treatment of a larger class of hyperparameters: RS does not require discretization of continuous spaces and also in principle allows also for distributions with non-compact supports.
The second row of Figure 3.1 shows uniform random samples with 4, 9 and 16 points on the unit square.
3.2.1 On the efficiency of random search (compared to grid search).
@@
In an influential article, bergstra_random_2012 advocate for using RS as the “default” baseline method for HPO instead of manual or grid search. They argue that while RS retains the advantages of GS, such as reproducibility, conceptual simplicity, very low implementation overhead and ease of parallelization, it is also more efficient in higher dimensional spaces, especially when the response function has a low effective dimensionality. The authors informally define this condition as when a subset of the hyperparameters accounts for most of the variation of the response function. For instance, for a bi-dimensional hyperparameter $\lambda=(\lambda_{1},\lambda_{2})$ it could be that $f(\lambda)\approx g(\lambda_{1})$, for some $g:\Lambda_{1}\to\mathbb{R}$.
@@
One way to reason about the success of RS is to assume that there is a sub-region $\Lambda^{*}\subset\Lambda$ that leads to a positive outcome of the experiment, i.e., where $f$ is sufficiently small. Then, we can assess the odds that a run of RS “lands” at least one configuration in the good region by comparing the volume of such region with that of the full search space under random sampling [brooks1958discussion]. We can model the probability of success of random search after $K$ draws as follows:
$\mathbb{P}_{(\lambda_{1},\dots\lambda_{K})\sim\upsilon^{K}}\left(\hat{\lambda}_{K}\in\Lambda^{*}\right)=1-\left(1-\mathbb{P}_{\lambda\sim\upsilon}(\lambda\in\Lambda_{*})\right)^{K},$ (3.3)
@@
where $\mathbb{P}_{\upsilon}$ denotes the probability of sampling from $\upsilon$, typically uniform or log-uniform. Eq. (3.3) says that the minimum number of trials needed to assure a positive outcome with a given probability, say $95\%$, depends only on the probability of hitting $\Lambda_{*}$ under $\upsilon$. This can be independent of the search space dimension if the response function has a low effective dimensionality. This argument may explain the success of RS as a baseline in seemingly complex settings such as neural architecture search [li2019random], where repeated experimentation on a few well-known benchmark datasets has led to the development of well-tuned search spaces for neural architectures.
@@
For a general search space, we should expect $\mathbb{P}_{\upsilon}(\lambda\in\Lambda_{*})$ to decay exponentially in some $m^{\prime}\leq m$. For concreteness, consider the case where $\Lambda=[0,1]^{m}$ is the $m$-dimensional unit hypercube and $\Lambda^{*}=[0,1/2]^{m^{\prime}}\times[0,1]^{m-m^{\prime}}$ is the “good region”. The $m^{\prime}$-dimensional 1/2-hypercube represents the important hyperparameters, while the $m-m^{\prime}$-dimensional unit hypercube represents the unimportant ones, as these coordinates do not matter in hitting $\Lambda^{*}$. If $\upsilon=\mathcal{U}(\Lambda)$ is the uniform distribution over $\Lambda$, then $\mathbb{P}_{\upsilon}(\lambda\in\Lambda_{*})=2^{-m^{\prime}}$. Consequently, the number of trials needed to obtain a positive outcome with at least $95\%$ probability is
$K\geq\log(0.05)/\log(1-2^{-m^{\prime}}).$
For $m^{\prime}=3$ one already needs $K\geq 23$ draws to achieve a $95\%$ probability of success. This illustrates that the number of trials needed grows exponentially in $m^{\prime}$. At the same time, this also shows that if $m^{\prime}\ll m$ RS is much more efficient compared to GS.
3.2.2 Unlucky runs
One potential disadvantage of RS is that its randomized nature means that “unlucky runs” may end up with entire regions of $\Lambda$ being under-explored or completely neglected. As a consequence, a single execution of random search may not necessarily provide sufficient information to meaningfully modify $\upsilon$ for another iteration of random search, should it be needed.
Let us illustrate this phenomenon. Continuing with our illustrative examples, say we divide the unit hypercube in $2^{m}$ 1/2 hypercubes $I_{j}$, for $j\in[2^{m}]$. We wish to have at least one configuration per hypercube (i.e., anywhere inside the 1/2 hypercube). While GS would need exactly $2^{m}$ points to achieve this, the probability under uniform sampling that RS hits all the hypercubes in $2^{m}$ draws is
$\mathbb{P}_{(\lambda_{1},\dots,\lambda_{2^{m}})\sim\upsilon^{2^{m}}}(\forall\,j\in[2^{m}],\;\;\exists\,k\in[2^{m}]\,\text{ s.t. }\,\lambda_{k}\in I_{j})=\prod_{j=1}^{2^{m}-1}(1-j/2^{m}).$
For example, for $m=3$ such probability is already $3/32$,. This means that, in expectation, RS with $8$ draws covers all the hypercubes only once every ten runs.
@@ 3.3 Quasi-random search
One basic requirement for a stochastic generator of hyperparameter configuration is consistency: we would like that $\lim_{k\to\infty}\min_{i\in[k]}f(\lambda_{i})=f(\lambda^{*})$. RS with a uniform distribution is consistent, while GS is not. However, unlike GS, random search for any fixed budget may generate a set of configurations that misses large portions of the search space. Intuitively, good sequences should cover evenly the search space. Also their projections onto subspaces of $\Lambda$ should be well spread, in order to speed up convergence when optimizing functions with low effective dimensionality – a property that GS misses. Quasi-random search methods seek to combine the advantages of grid and random search, with configuration-drawing strategies that combine randomness with deterministic steps.
Formalizations of these desiderata have been developed in the context of numerical integration and global optimization via the concepts of low dispersion and low discrepancy sequences [see e.g. sobol1979systematic, niederreiter1988low, dick2010digital]. bousquet2017critical study the usage of such quasi-random sequences in HPO, finding that they consistently outperform random search (with a uniform distribution) on a series of deep learning tasks; see also [tuningplaybookgithub, Chapter 7] The resulting methods are conceptually very similar to both GS and RS, being highly parallelizable, except that the set of hyperparameters $\lambda_{1},\dots,\lambda_{k}$ are generated according to more complex procedures. We review two possible choices below, assuming that the search space is an $m$-dimensional unit hypercube for simplicity. Note that this is typically not at all an hard limitation: starting with samples on the unit interval, one can easily derive samples for a variety of distributions with appropriate reparameterizations [see, e.g. devroye2006nonuniform].
3.3.1 Latin hypercube sampling (LHS)
LHS [mckay1979comparison] directly combines grid and random search, in that it divides the search space into a grid, but samples configurations at random within regions of the grid.
LHS works as follows: choose a sequence length $n>1$, which determines the number of hyperparameters generated by a run of the sampling process. Divide each $[0,1]$ interval of the unit hypercube into $n>1$ parts, denoting by $I_{s}=[(s-1)/n,s/n]$ the resulting intervals. This gives rise to $n^{m}$ hypercubes of volume $n^{-m}$ each. Then generate $m$ random permutations of $[n]$, denoted by $\{\sigma_{i}\}_{i=1}^{m}$, $\sigma_{i}:[n]\to[n]$. For $k\in[n]$, sample hyperparameter configurations as follows:
@@ 3.3.1 Latin hypercube sampling (LHS)
where $\mathcal{U}(H_{k})$ is the uniform distribution on $H_{k}$. This generates $n$ trial points; one may then generate further sequences of $n$ configurations by repeating the process. Importantly, Latin hypercube sampling ensures that the projection on each axis of the $n$-long sequences places exactly one sample per interval.
The second last row of Figure 3.1 shows Latin hypercube samples on the unit square with 4, 9 and 16 points; namely, for $n\in\{2,3,4\}$.
3.3.2 Sobol’ sequences
Sobol’ sequences [sobol1979systematic] are a collection of separate one-dimensional sequences, one for each axis: when constructing a $m+m^{\prime}$ dimensional sequence one can reuse the $m$-dimensional one for the first coordinates.
Let $j\in[m]$ be an axis, and let $\mathbb{F}_{2}=\{0,1\}$ with integer operations modulo $2$. For each integer $k$ – representing the iterate – let $k=\sum_{l}a_{l}(k)2^{j}$ be the binary expansion of $k$, where $a_{l}(k)\in\mathbb{F}_{2}$ are binary coefficients (e.g., for $k=3$, these would be $a_{0}(3)=1$, $a_{1}(3)=1$ and the rest 00’s). Denote by $a(k)$ the resulting vector. Let $C_{j}$ be a specific binary (infinite) “direction matrix” for axis $j$. Then the $j$-th coordinate of the $k$-th iterate of a Sobol’ sequence can be written as
@@ 3.3.2 Sobol’ sequences
where the matrix-vector multiplication $C^{j}\cdot a(k)$ is in $\mathbb{F}_{2}$ (i.e., the additions are modulo 2). The direction matrices $C^{j}$ are constructed starting from primitive polynomials of $\mathbb{F}_{2}$ and direction numbers, which must be carefully chosen to ensure good coverage of the sequences. We refer the reader to [Lemieux:1338331, Chapter 5] for details.
Note that, given the directions matrices, Sobol’ sequences are deterministic. It is common practice to obtain stochastic variants of such sequences by adding random shifts, i.e. $\lambda^{\prime}_{k}=\lambda_{k}+u$ (modulo the unit hypercube) with $u\sim\mathcal{U}([0,1]$ and by “scrambling” the binary representation vector $a(k)$, that is, randomly permuting its entries [owen1995randomly].
The bottom row of Figure 3.1 plots 4, 9 and 16 points Sobol sequences with random shift and scrambling on the unit square.
3.3.3 Other schemes
@@
Figure 3.1: Non-adaptive sampling methods on the unit square for different sample sizes: 4 (left), 9 (center) and 16 (right) points. Each row represents a different schema. Grid and Latin hypercube sampling (LHS) may require re-sampling every time we increase the sample size. Instead, for random and Sobol we can keep on adding points, “reusing” previous samples. For each plot, we also show projections of the samples onto the two axes. We can see how grid samples cover much fewer points on the projections w.r.t. the other sampling methods. We can also observe the “unlucky run” phenomenon of (uniform) random sampling, where after nine draws the upper-right quadrant is still unvisited. LHS and Sobol (with random shifts and scrambling) provide more even coverage of the space and its unidimensional projections.
@@
The techniques described in the previous chapter do not make any attempt to model the response function $f(\lambda)$: in other words, evaluations of $f$ on previously selected configurations do not influence the search. In contrast, model-based (MB) techniques construct an explicit surrogate model for $f(\lambda)$ – or for closely related quantities – based on the information collected from previous evaluations. The surrogate model is a first central component of model-based HPO methods. We will denote such model by
@@
The landscape of model-based methods is strongly dominated by techniques falling under the umbrella of Bayesian optimization (BO), which we cover in Section 4.2. In BO, the surrogate model is a probabilistic mapping that also captures the uncertainty of the estimation; implementations include Gaussian processes (Section 4.2.2) and tree Parzen estimators (Section 4.2.5). Regardless of the particular implementation, one critical feature of surrogate models is that they should be much cheaper to evaluate than $f$ itself. BO methods do not locally minimize the surrogate function to select the next configuration to evaluate, but rather mediate it through a so-called acquisition function (AF), a second central component of a BO method and defined as a mapping:
$\alpha:\Lambda\times\hat{\mathcal{F}}\times\mathcal{X}\to\mathbb{R}$ (4.2)
where $\hat{\mathcal{F}}$ is an appropriate space of surrogate models, and, again, $\mathcal{X}$ collects auxiliary inputs, that is used to select the next configuration to evaluate $f$ on. Typically, $\alpha$ encodes a trade-off between exploration and exploitation:
  • exploration seeks to evaluate $f$ on points with high uncertainty, with the primary aim to improve the surrogate model;
  • exploitation seeks to leverage knowledge gathered so far to improve on the incumbent solution.
We discuss acquisition functions in Section 4.2.3.
We briefly touch upon other approaches to model-based optimization that do not directly fall in the BO category in Section 4.3. We conclude the chapter with a practical example that contrasts BO to random search 4.4 and an overall discussion in Section 4.5. In particular, we discuss parallelizability of model-based methods, a challenging practical bottleneck of MB-HPO methods (Section 4.5.1).
4.1 Model-based HPO: a general schema
Algorithm 1 Model-based-HPO
1:Response function $f$
2:Warm-up dataset $\mathcal{C}_{0}=\{(\lambda_{j},f(\lambda_{j}))\}$
3:Class of surrogate functions $\hat{\mathcal{F}}$ and fitting routine
4:Acquisition function $\alpha$
5:$\beta\leftarrow\min_{j}f(\lambda_{j})$triangleright Initialize incumbent
6:$\mathcal{C}_{1}\leftarrow\mathcal{C}_{0}$
7:for $k\leftarrow 1$ to $T$ do
8:     Fit $\hat{f}_{k}$ to $\mathcal{C}_{k}$triangleright Update step ($\pi_{2}$)
9:     $\lambda_{k}\leftarrow\operatorname*{arg\,max}_{\lambda^{\prime}}\alpha(\lambda^{\prime}|\hat{f}_{k-1},\beta)$triangleright Decision step ($\pi_{1}$)
10:     Evaluate $f(\lambda_{k})$triangleright Expensive step
11:     if $f(\lambda_{k})\leq\beta)$ then
12:         $\beta\leftarrow f(\lambda_{k})$
13:     end if
14:     $\mathcal{C}_{k+1}\leftarrow\mathcal{C}_{k}\cup(\lambda,f(\lambda))$
15:end for